OpenStudy (anonymous):

Solve 9[8 - (8x - 6)] + 3x = 0.

HanAkoSolo (jamierox4ev3r):

hello @NikkiToh , the key to solving problems like these is to distribute

HanAkoSolo (jamierox4ev3r):

We are trying to solve for the variable x, correct?

HanAkoSolo (jamierox4ev3r):



OpenStudy (anonymous):

Actually @Jamierox4ev3r I don't know how to solve that.

OpenStudy (anonymous):

That is my teacher question.

HanAkoSolo (jamierox4ev3r):

oh okay. So try to distribute within the parentheses first. for example, -(8x-6) would become -8x+6 Then, you would have 9(8-(-8x+6))+3x=0 Do you understand what I've done so far? @NikkiToh

OpenStudy (anonymous):

Okay @Jamierox4ev3r I will follow you, but what is the final answer in that question ?

OpenStudy (anonymous):

@Jamierox4ev3r its already done, but thankyou for your time. :*

HanAkoSolo (jamierox4ev3r):

hold on, let me explain first. once you have 9(8-(-8x+6))+3x=0, distribute again Then you can simplify it to this equation: 126-72x+3x=0 oh lol okay, what answer did you submit?

OpenStudy (godlovesme):

Welcome to OS :)) @nikkitoh and nice nob jamie :)

HanAkoSolo (jamierox4ev3r):

hey thanks @Godlovesme I'm kinda half asleep right now, I hate it when the person I'm trying to help just leaves...but lol it's okay i guess :)

OpenStudy (godlovesme):

Lol me too :P but its alright and yw =)

OpenStudy (anonymous):

Im sorry for that Jaime. Because were already done for that subject. But I would like to say thankyou for your help. Don't be mad at me please. :)